if product demand is given by: q(p)= 5000-100 p, what is the marginal revenue function for this product
Answer #1

Ann. We have q (p) = 5000 - 100 (p) CP) = 5000 - qucp) loo = so - 1 q cp) 10o As, Total Revenue - = P x (a) px Corp) So TR =